OpenStudy (anonymous):

A right circular cylinder has a total surface area of 112pie. If the height of the cylinder is 10, find the diameter of the base.

OpenStudy (anonymous):

haha you called it 112"pie". that's funny. formula is circumference * height + 2 (area of circle) that works out to 2pi*r*h + 2pi*r^2 = 112pi You have h so you can plug that in. 20pi*r + 2pi*r^2 = 112pi cancel out the pi. 20r + 2r^2=112 move stuff over. 2r^2 + 20r - 112 = 0 divide. r^2 + 10r - 56 = 0 factor. (r+14)(r-4) -14 and 4 are the answers, only positive number works, so r = 4 d = 8

OpenStudy (anonymous):

hey i had answered u already diameter is approximately equal to 83 mm and exactly it is 82.8 mm

OpenStudy (anonymous):

I got that one Sheg but it was for a different question. Thank You!